# Lattice determination of the QCD low-energy constant $\ell_{\scriptscriptstyle{7}}$

### v1

@@ Abstract
We provide a non-perturbative determination of the scheme- and scale-independent low-energy constant $\ell_{\scriptscriptstyle{7}}$, appearing in the QCD effective chiral Lagrangian at next-to-leading order, by means of lattice QCD simulations with $N_{\scriptscriptstyle{\rm f}}=2+1$ quark flavors. We adopt staggered fermions and extract $\ell_{\scriptscriptstyle{7}}$ from the pion mass splitting by suitably generalizing the method introduced in [Phys. Rev. D 104 (2021) 074513] for the Wilson discretization. Adopting 12 gauge ensembles with 3 different values of the pion mass, and 4 different values of the lattice spacing, we are able to achieve controlled extrapolations towards the continuum, infinite volume, and chiral limits. Our final result $\ell_{\scriptscriptstyle{7}}\,\times\,10^{3}=1.98(48)_{\scriptscriptstyle{\rm stat}}(26)_{\scriptscriptstyle{\rm syst}}=1.98(54)_{\scriptscriptstyle{\rm tot}}$ agrees with and substantially improves on previous determinations.
@@ II.1 Lattice QCD discretization and simulation parameters
$0.1421$ 3.678 20 0.0088 0.0621 0.1515 6400 260(3) 112.7(1.1) 3.03 4.00
3.750 24 0.0072 0.0503 0.1265 5600 262(3) 115.5(1.3) 3.04 4.08
3.868 32 0.0054 0.0379 0.0964 3573 263(4) 116.6(1.2) 3.08 4.21
3.988 40 0.0042 0.0299 0.0758 3360 269(5) 114.7(1.2) 3.03 4.13
$0.2131$ 3.678 20 0.0132 0.0621 0.1532 2000 320(2) 118.1(0.9) 3.06 4.96
3.750 24 0.0107 0.0503 0.1278 2000 315(3) 121.5(0.7) 3.07 4.90
3.868 32 0.0081 0.0379 0.0976 2930 315(5) 122.3(1.1) 3.12 4.98
3.988 40 0.0064 0.0299 0.0764 1590 317(6) 122.1(0.9) 3.06 4.92
$0.3197$ 3.678 20 0.0199 0.0621 0.1556 1520 381(1) 125.4(0.6) 3.11 6.00
3.750 24 0.0161 0.0503 0.1297 1480 382(2) 128.1(0.6) 3.11 6.02
3.868 32 0.0121 0.0379 0.0989 1450 380(1) 129.4(0.7) 3.16 6.09
3.988 40 0.0095 0.0299 0.0768 1520 386(4) 128.3(0.6) 3.07 6.01
Table 1: Summary of the simulation parameters of the gauge ensembles employed in this study. The LCP parameters, including pion masses and lattice spacings, were obtained in [15] starting from the physical point LCP determined in [16, 17, 18]. The values of the lattice spacing have been determined in [15] with the $w_{0}$ scale setting approach [19] with a $\sim 2\%$ error, and their value in fm units was obtained assuming $w_{0}=0.1757(12)~\mathrm{fm}$ [19] for all ensembles, independently of the pion mass. The variable $N_{\rm conf}$ is the number of thermalized gauge configurations generated for each ensemble, each separated by 30 RHMC updating steps. For each ensemble we also report the values of $M_{\pi}$ and $F_{\pi}$, obtained in [15].
@@
In this work we will employ 12 different gauge ensembles, whose simulation parameters are reported in Tab. 1. These ensembles correspond to 3 different Lines of Constant Physics (LCPs) where the lattice size is fixed to $L\equiv aN_{\scriptscriptstyle{\rm s}}\simeq 3~\mathrm{fm}$, the bare strange quark mass $m_{\scriptscriptstyle{\rm s}}$ is tuned to keep its renormalized value constant and equal to its physical value $m_{\scriptscriptstyle{\rm s}}^{{\scriptscriptstyle{(\mathrm{phys})}}}$, and the bare mass of the degenerate light doublet $m_{\scriptscriptstyle{\rm u}}=m_{\scriptscriptstyle{\rm d}}=m_{\scriptscriptstyle{\ell}}$ is varied to explore different values of the pion mass $M_{\pi}$. These LCPs were determined in Ref. [15], starting from the results of Refs. [16, 17, 18] for the bare parameters belonging to the physical point LCP, by varying the bare light quark mass at fixed bare strange quark mass. More precisely, we considered light-to-strange quark mass ratios $R\equiv m_{\scriptscriptstyle{\ell}}/m_{\scriptscriptstyle{\rm s}}^{{\scriptscriptstyle{(\mathrm{phys})}}}$ equal to 4, 6 and 9 times the physical ratio $R^{{\scriptscriptstyle{(\mathrm{phys})}}}=m_{\scriptscriptstyle{\ell}}^{{\scriptscriptstyle{(\mathrm{phys})}}}/m_{\scriptscriptstyle{\rm s}}^{{\scriptscriptstyle{(\mathrm{phys})}}}$ determined in Refs. [16, 17, 18]. These choices correspond to $M_{\pi}^{2}\simeq 4,~6,~9\,{M_{\pi}^{{\scriptscriptstyle{(\mathrm{phys})}}}}^{2}$, $M_{\pi}^{{\scriptscriptstyle{(\mathrm{phys})}}}=135$ MeV, and $M_{\pi}L\geq 4$, a value which is sufficient to keep finite-size effects below our statistical errors, as we shall discuss later. For each LCP we considered 4 values of the lattice spacing $a$ ranging from $\sim 0.15$ fm to $\sim 0.075$ fm. Our setup thus allows us to take continuum limits $a\to 0$ at fixed light and strange quark masses, and then to take the SU(2) chiral limit $m_{\scriptscriptstyle{\ell}}\to 0$ at fixed physical strange quark mass $m_{\scriptscriptstyle{\rm s}}^{{\scriptscriptstyle{(\mathrm{phys})}}}$, i.e., $R\to 0$. Further details about the determination of the LCP simulation parameters can be found in the original paper [15].
@@ II.2 $\ell_{\scriptscriptstyle{7}}$ from the mass-splitting method
The starting point is to rewrite (1) in a way which is more suitable for a numerical evaluation. By expanding the left-hand side at LO in $\Delta m$ and using the $\chi\mathrm{PT}$ LO relation $M^{2}=B(m_{\scriptscriptstyle{\rm u}}+m_{\scriptscriptstyle{\rm d}})=2Bm_{\scriptscriptstyle{\ell}}$, we have:
@@ III.1 Lattice determination of $\ell_{\scriptscriptstyle{7}}$
Instead, concerning the 4-point correlators $C_{{\scriptscriptstyle{4\mathrm{pt}}}}^{{\scriptscriptstyle{(\rm conn)}}}$ and $C_{{\scriptscriptstyle{4\mathrm{pt}}}}^{{\scriptscriptstyle{(\rm disc)}}}$, we found that the standard pseudoscalar channel revealed to be impractical, as with this choice of the pion interpolating operator the disconnected correlator turned out to be completely dominated by statistical noise, and remained compatible with zero within errors for all time separations [^1] A similar behavior for disconnected diagrams in the pseudoscalar channel was observed in Ref. [26]. In that study it is pointed out that a term appearing in the effective Staggered $\chi\mathrm{PT}$ Lagrangian allows disconnected contributions in the meson propagator only in the singlet, vector, and axial vector channels. Thus, to study the $\eta$ and $\eta^{\prime}$ mesons, where disconnected diagrams are crucial due to their flavor-singlet nature, the $\gamma_{5}\times\mathrm{I}$ interpolating operator was employed, as it is the only pseudoscalar bilinear with a non-vanishing disconnected contribution in the staggered case.. Thus, the pseudoscalar channel does not allow for a reliable extraction of $\ell_{\scriptscriptstyle{7}}$, motivating the need to explore alternative spin-taste structures. The first natural candidate beyond the pseudoscalar channel is the axial-vector one, corresponding to $\xi_{{\scriptscriptstyle{\rm A}}}=\xi_{5\mu}$. In this case, both the connected and disconnected diagrams exhibit non-zero signal with reasonable statistical quality, thus allowing a reliable calculation of the estimator $\ell_{\scriptscriptstyle{7}}^{{\scriptscriptstyle{(\rm eff)}}}(t)$ (see Appendix B for a comparison of the obtained results for the disconnected contribution to $\ell_{\scriptscriptstyle{7}}$ from the pseudo-scalar and the axial channels). As a further check, we also verified that compatible results for $\ell_{\scriptscriptstyle{7}}$ are obtained from the vector channel ($\xi_{{\scriptscriptstyle{\rm V}}}=\xi_{\mu}$) once the continuum limit is taken, although within larger error bars, see Appendix C, thus motivating our choice of the axial-vector channel.
Figure 1: Left panel: lattice estimator $\ell_{\scriptscriptstyle{7}}^{{\scriptscriptstyle{(\rm eff)}}}(t)$ and individual contributions of the connected/disconnected diagrams to this function for the ensemble with $R\simeq 0.1421$ and $a=0.095$ fm. Right panel: extraction of $\ell_{\scriptscriptstyle{7}}$ from the large-time-separation plateau exhibited by $\ell_{\scriptscriptstyle{7}}^{{\scriptscriptstyle{(\rm eff)}}}(t)$.
@@
We conclude by reporting for comparison the result from the mass splitting method obtained in Ref. [11], for a lattice spacing $a\simeq 0.095$ fm, a pion mass $M_{\pi}\simeq 260$ MeV, and a volume $M_{\pi}L\simeq 4$:
@@
It should be noted that the error reduction achieved here cannot be entirely ascribed to the different sample sizes, as our statistics is larger by about a factor of 3 with respect to the one of [11], but we find a smaller uncertainty by about a factor of 5.
Figure 2: Study of finite-size effects on $\ell_{\scriptscriptstyle{7}}$ for our ensemble with $a=0.0964$ $\mathrm{fm}$ and $M_{\pi}=263(4)$ $\mathrm{MeV}$. They are invisible within our statistical errors for $M_{\pi}L\geq 4$. Same conclusions are reached after subtracting finite-size effects estimated from NLO $\chi\mathrm{PT}$: $\ell_{\scriptscriptstyle{7}}(L)\to\ell_{\scriptscriptstyle{7}}(L)/[1+R_{\ell_{\scriptscriptstyle{7}}}(L)]$, see Sec. III.2.
@@ III.2 Finite-size effects on $\ell_{\scriptscriptstyle{7}}$
For the next-to-finest lattice spacing of our lightest ensembles with $M_{\pi}L\simeq 4$ we explicitly checked that this threshold is sufficient to keep finite-size effects under control by performing two additional simulations for a smaller and a larger value of $M_{\pi}L\simeq 3,5$. The result of this test, shown in Fig. 2, indeed confirm that our choice for the lattice volume is safe. Such conclusion is also supported by analytical results obtained from NLO $\chi\mathrm{PT}$. Indeed, analytic expressions for the volume-dependence of $M_{\pi}$ and $F_{\pi}$ are available up to NLO in the chiral expansion [27, 28] (for the pion mass, also NNLO effects are known [29] and have been included in our calculation):
@@ III.3 Continuum extrapolations of $\ell_{\scriptscriptstyle{7}}$
Figure 3: Continuum limit extrapolations of $\ell_{\scriptscriptstyle{7}}$ for the 3 values of the parameter $R=m_{\scriptscriptstyle{\ell}}/m_{\scriptscriptstyle{\rm s}}^{{\scriptscriptstyle{(\mathrm{phys})}}}$ explored. Left panels report best fits as a function of $a^{2}$, while right panels report best fits as a function of $w_{0}^{2}\Delta$. Continuous lines (with their related continuous shaded areas) correspond to linear fits using all available data points, dashed lines represent linear fits restricted to the 3 finest lattice spacings, and dotted lines (with their related dashed shaded areas) represent quadratic fits. The star points represent our final results for the continuum limits, are obtained from the linear extrapolations in $a^{2}$ (no star point is represented for extrapolations in $w_{0}^{2}\Delta$, as these fits are only used to assess systematic errors). Since the total error bar, given by the sum in quadrature of the statistical and systematic uncertainties, is barely distinguishable from the statistical one for the two largest values of $R$, it has been slightly shifted horizontally for clarity.
@@
Given the overall good agreement among all possible continuum limit fit functions, in all cases we considered the outcome of the linear fit in $a^{2}$, performed considering all lattice spacings, to assign a central value and a statistical error to our continuum determinations of $\ell_{\scriptscriptstyle{7}}(R)$.
In addition to statistical uncertainties, we also assigned to our continuum extrapolations a systematic error related to the small difference observed when extrapolating our data with different fit Ansätze. Following the procedure outlined in [33, 15], we computed:

### v2

@@ Abstract
We provide a non-perturbative determination of the scheme- and scale-independent low-energy constant $\ell_{\scriptscriptstyle{7}}$, appearing in the QCD effective chiral Lagrangian at next-to-leading order, by means of lattice QCD simulations with $N_{\scriptscriptstyle{\rm f}}=2+1$ quark flavors. We adopt staggered fermions and extract $\ell_{\scriptscriptstyle{7}}$ from the pion mass splitting by suitably generalizing the method introduced in [Phys. Rev. D 104 (2021) 074513] for the Wilson discretization. Adopting 12 gauge ensembles with 3 different values of the pion mass, and 4 different values of the lattice spacing, we are able to achieve controlled extrapolations towards the continuum, infinite volume, and chiral limits. Our final result $\ell_{\scriptscriptstyle{7}}\,\times\,10^{3}=2.79(58)_{\scriptscriptstyle{\rm stat}}(19)_{\scriptscriptstyle{\rm syst}}=2.79(61)_{\scriptscriptstyle{\rm tot}}$ agrees with and substantially improves on previous determinations.
@@ II.1 Lattice QCD discretization and simulation parameters
$0.1421$ 3.67838 20 0.008828 0.062136 0.1515 6200 260(3) 98.5(2.1) 3.03 4.00
3.75000 24 0.007150 0.050300 0.1265 5400 262(3) 99.0(2.3) 3.04 4.08
3.86847 32 0.005383 0.037884 0.0964 3373 263(4) 98.9(2.2) 3.08 4.21
3.98775 40 0.004244 0.029867 0.0758 3160 269(5) 98.4(2.3) 3.03 4.13
$0.2131$ 3.67838 20 0.013242 0.062136 0.1532 1800 320(2) 102.4(2.7) 3.06 4.96
3.75000 24 0.010722 0.050300 0.1278 1800 315(3) 101.8(2.5) 3.07 4.90
3.86847 32 0.008075 0.037884 0.0976 2730 315(5) 100.9(2.4) 3.12 4.98
3.98775 40 0.006366 0.029867 0.0764 1390 317(6) 101.1(2.5) 3.06 4.92
$0.3197$ 3.67838 20 0.019863 0.062136 0.1556 1320 381(1) 106.8(2.6) 3.11 6.00
3.75000 24 0.016080 0.050300 0.1297 1280 382(2) 105.8(2.4) 3.11 6.02
3.86847 32 0.012112 0.037884 0.0989 1250 380(1) 105.5(2.5) 3.16 6.09
3.98775 40 0.009549 0.029867 0.0768 1320 386(4) 106.3(2.6) 3.07 6.01
Table 1: Summary of the simulation parameters of the gauge ensembles employed in this study. The LCP parameters, including pion masses and lattice spacings, were obtained in [15] starting from the physical point LCP determined in [16, 17, 18]. The values of the lattice spacing have been determined in [15] with the $w_{0}$ scale setting approach [19] with a $\sim 2\%$ error at most, and their value in fm units was obtained assuming $w_{0}=0.1757(12)~\mathrm{fm}$ [19] for all ensembles, independently of the pion mass. The variable $N_{\rm conf}$ is the number of thermalized gauge configurations generated for each ensemble, each separated by 10 RHMC updating steps. For each ensemble we also report the values of $M_{\pi}$ and $F_{\pi}$.
@@
In this work we will employ 12 different gauge ensembles, whose simulation parameters are reported in Tab. 1. These ensembles correspond to 3 different Lines of Constant Physics (LCPs) where the lattice size is fixed to $L\equiv aN_{\scriptscriptstyle{\rm s}}\simeq 3~\mathrm{fm}$, the bare strange quark mass $m_{\scriptscriptstyle{\rm s}}$ is tuned to keep its renormalized value constant and equal to its physical value $m_{\scriptscriptstyle{\rm s}}^{{\scriptscriptstyle{(\mathrm{phys})}}}$, and the bare mass of the degenerate light doublet $m_{\scriptscriptstyle{\rm u}}=m_{\scriptscriptstyle{\rm d}}=m_{\scriptscriptstyle{\ell}}$ is varied to explore different values of the pion mass $M_{\pi}$. These LCPs were determined in Ref. [15], starting from the results of Refs. [16, 17, 18] for the bare parameters belonging to the physical point LCP, by varying the bare light quark mass at fixed bare strange quark mass. More precisely, we considered light-to-strange quark mass ratios $R\equiv m_{\scriptscriptstyle{\ell}}/m_{\scriptscriptstyle{\rm s}}^{{\scriptscriptstyle{(\mathrm{phys})}}}$ equal to 4, 6 and 9 times the physical ratio $R^{{\scriptscriptstyle{(\mathrm{phys})}}}=m_{\scriptscriptstyle{\ell}}^{{\scriptscriptstyle{(\mathrm{phys})}}}/m_{\scriptscriptstyle{\rm s}}^{{\scriptscriptstyle{(\mathrm{phys})}}}$ determined in Refs. [16, 17, 18]. These choices correspond to $M_{\pi}^{2}\simeq 4,~6,~9\,{M_{\pi}^{{\scriptscriptstyle{(\mathrm{phys})}}}}^{2}$, $M_{\pi}^{{\scriptscriptstyle{(\mathrm{phys})}}}=135$ MeV, and $M_{\pi}L\geq 4$, a value that is sufficient to keep finite-size effects below our statistical errors, as we shall discuss later. For each LCP we considered 4 values of the lattice spacing $a$ ranging from $\sim 0.15$ fm to $\sim 0.075$ fm. Our setup thus allows us to take continuum limits $a\to 0$ at fixed light and strange quark masses, and then to take the SU(2) chiral limit $m_{\scriptscriptstyle{\ell}}\to 0$ at fixed physical strange quark mass $m_{\scriptscriptstyle{\rm s}}^{{\scriptscriptstyle{(\mathrm{phys})}}}$, i.e., $R\to 0$. Further details about the determination of the LCP simulation parameters can be found in the original paper [15]. [^1] The results for $F_{\pi}$ in Tab. 1 correct those originally given in Ref. [15], whose calculation was affected by a mismatch in the renormalization constants, resulting in a $\sim 15-19\%$ shift upward. This issue regarded the analysis of the ensembles at heavier-than physical pions, and it does not affect the chiral limit of $F_{\pi}$ given in [15], but only its slope as a function of $M_{\pi}$. The latter was not discussed in [15], and it is reported for the first time in the present work in Appendix E.
@@ II.2 $\ell_{\scriptscriptstyle{7}}$ from the mass-splitting method
The starting point is to rewrite (1) in a way that is more suitable for a numerical evaluation. By expanding the left-hand side at LO in $\Delta m$ and using the $\chi\mathrm{PT}$ LO relation $M^{2}=B(m_{\scriptscriptstyle{\rm u}}+m_{\scriptscriptstyle{\rm d}})=2Bm_{\scriptscriptstyle{\ell}}$, we have:
@@ III.1 Lattice determination of $\ell_{\scriptscriptstyle{7}}$
Figure 1: Left panel: lattice estimator $\ell_{\scriptscriptstyle{7}}^{{\scriptscriptstyle{(\rm eff)}}}(t)$ and individual contributions of the connected/disconnected diagrams to this function for the ensemble with $R\simeq 0.1421$ and $a=0.0964$ fm. Right panel: extraction of $\ell_{\scriptscriptstyle{7}}$ from the large-time-separation plateau exhibited by $\ell_{\scriptscriptstyle{7}}^{{\scriptscriptstyle{(\rm eff)}}}(t)$.
Instead, concerning the 4-point correlators $C_{{\scriptscriptstyle{4\mathrm{pt}}}}^{{\scriptscriptstyle{(\rm conn)}}}$ and $C_{{\scriptscriptstyle{4\mathrm{pt}}}}^{{\scriptscriptstyle{(\rm disc)}}}$, we found that the standard pseudoscalar channel revealed to be impractical, as with this choice of the pion interpolating operator the disconnected correlator turned out to be completely dominated by statistical noise, and remained compatible with zero within errors for all time separations [^2] A similar behavior for disconnected diagrams in the pseudoscalar channel was observed in Ref. [26]. In that study it is pointed out that a term appearing in the effective Staggered $\chi\mathrm{PT}$ Lagrangian allows disconnected contributions in the meson propagator only in the singlet, vector, and axial vector channels. Thus, to study the $\eta$ and $\eta^{\prime}$ mesons, where disconnected diagrams are crucial due to their flavor-singlet nature, the $\gamma_{5}\times\mathrm{I}$ interpolating operator was employed, as it is the only pseudoscalar bilinear with a non-vanishing disconnected contribution in the staggered case.. Thus, the pseudoscalar channel does not allow for a reliable extraction of $\ell_{\scriptscriptstyle{7}}$, motivating the need to explore alternative spin-taste structures. The first natural candidate beyond the pseudoscalar channel is the axial-vector one, corresponding to $\xi_{{\scriptscriptstyle{\rm A}}}=\xi_{5\mu}$. In this case, both the connected and disconnected diagrams exhibit non-zero signal with reasonable statistical quality, thus allowing a reliable calculation of the estimator $\ell_{\scriptscriptstyle{7}}^{{\scriptscriptstyle{(\rm eff)}}}(t)$ (see Appendix B for a comparison of the obtained results for the disconnected contribution to $\ell_{\scriptscriptstyle{7}}$ from the pseudo-scalar and the axial channels). As a further check, we also verified that compatible results for $\ell_{\scriptscriptstyle{7}}$ are obtained from the vector channel ($\xi_{{\scriptscriptstyle{\rm V}}}=\xi_{\mu}$) once the continuum limit is taken, although within larger error bars, see Appendix B, thus motivating our choice of the axial-vector channel.
@@
We conclude by reporting for comparison the result from the mass splitting method obtained in Ref. [11] with twisted mass Wilson fermions, for a lattice spacing $a\simeq 0.095$ fm, a pion mass $M_{\pi}\simeq 260$ MeV, and a volume $M_{\pi}L\simeq 4$:
@@
It should be noted that the error reduction achieved here cannot be entirely ascribed to the different sample sizes, as our statistics is larger by about a factor of 3 with respect to the one of [11], but we find a smaller uncertainty by about a factor of 5.  [^3] On the other hand, the statistical accuracies of $F_{\pi}$ and $M_{\pi}$ obtained with staggered quarks here and with twisted mass Wilson fermions in [11] are similar. More comprehensive investigations with other discretizations and possibly with mixed action setups are needed to clarify how the statistical accuracies of different types of observables depend on the chosen fermion formulation.
@@ III.2 Finite-size effects on $\ell_{\scriptscriptstyle{7}}$
For the next-to-finest lattice spacing of our lightest ensembles with $M_{\pi}L\simeq 4$ we explicitly checked that this threshold is sufficient to keep finite-size effects under control by performing two additional simulations for a smaller and a larger value of $M_{\pi}L\simeq 3$ and $5$. The result of this test, shown in Fig. 2, indeed confirm that our choice for the lattice volume is safe. Such conclusion is also supported by analytical results obtained from NLO $\chi\mathrm{PT}$. Indeed, analytic expressions for the volume-dependence of $M_{\pi}$ and $F_{\pi}$ are available up to NLO in the chiral expansion [27, 28] (for the pion mass, also NNLO effects are known [29] and have been included in our calculation):
@@
Figure 2: Study of finite-size effects on $\ell_{\scriptscriptstyle{7}}$ for our ensemble with $a=0.0964$ $\mathrm{fm}$ and $M_{\pi}=263(4)$ $\mathrm{MeV}$. They are invisible within our statistical errors for $M_{\pi}L\geq 4$. The same conclusions are reached after subtracting finite-size effects estimated from NLO $\chi\mathrm{PT}$: $\ell_{\scriptscriptstyle{7}}(L)\to\ell_{\scriptscriptstyle{7}}(L)/[1+R_{\ell_{\scriptscriptstyle{7}}}(L)]$, see Sec. III.2.
@@ III.3 Continuum extrapolations of $\ell_{\scriptscriptstyle{7}}$
Figure 3: Continuum limit extrapolations of $\ell_{\scriptscriptstyle{7}}$ for the 3 values of the parameter $R=m_{\scriptscriptstyle{\ell}}/m_{\scriptscriptstyle{\rm s}}^{{\scriptscriptstyle{(\mathrm{phys})}}}$ explored. Left panels report best fits as a function of $a^{2}$, while right panels report best fits as a function of $w_{0}^{2}\Delta$. Continuous lines (with their related continuous shaded areas) correspond to linear fits using all available data points, dashed lines represent linear fits restricted to the 3 finest lattice spacings, and dotted lines represent quadratic fits. The star points represent our final results for the continuum limits, and are obtained from the linear extrapolations in $a^{2}$ (no star point is represented for extrapolations in $w_{0}^{2}\Delta$, as these fits are only used to assess systematic errors). Since the total error bar, given by the sum in quadrature of the statistical and systematic uncertainties, is barely distinguishable from the statistical one for the two largest values of $R$, it has been slightly shifted horizontally for clarity.
@@
Given the overall good agreement among all possible continuum limit fit functions, in all cases we considered the outcome of the linear fit in $a^{2}$, performed considering all lattice spacings, to assign a central value and a statistical error to our continuum determinations of $\ell_{\scriptscriptstyle{7}}(R)$. In addition to statistical uncertainties, we also assigned to our continuum extrapolations a systematic error related to the small difference observed when extrapolating our data with different fit Ansätze. Following the procedure outlined in [33, 15], we computed: